# DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition

### v4

@@ Appendix C Detail of DynamicNER
Appendix D More Experiment about CascadeNER
D.1 CascadeNER Setting
In the experiments of this section, CascadeNER always employs two Qwen2.5-7B base models, which are fine-tuned separately based on the corresponding part of the dynamic version of DynamicNER to obtain an extractor and a classifier. Potential data contamination about the fine-tuning is discussed in Appendix H. We evaluate CascadeNER’s performance in both few-shot and zero-shot scenarios, comparing it with supervised SOTAs and LLM-based baselines. For few-shot scenarios, the number of few-shot demonstrations is set to $3$, the same as the experiments on DynamicNER.
D.2 Baselines
@@
D.3 Dataset
Few-shot Data Sampling
Dataset Selection
@@ D.3 Dataset
CoNLL2003
@@
MultiCoNER & PAN-X
MultiCoNER and PAN-X are two widely used multilingual datasets. MultiCoNER covers 6 entity types and 11 languages, while PAN-X includes 3 entity types and 282 languages. We use 6 and 8 overlapping languages from MultiCoNER and PAN-X with DynamicNER to evaluate CascadeNER’s multilingual capabilities. It is important to note that, for the purpose of controlling variables, all methods requiring training are trained using multilingual joint training.
D.4 Experimental Results
As shown in Table 6 and 7, the results indicate that in low-resource scenarios, LLM-based methods achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music and FewNERD, and PAN-X and MultiCoNER in some languages, achieving new SOTA performance and highlighting its exceptional generalization and capability to handle complex entity categorization. However, when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X, indicating that supervised methods are still useful in some scenerios.
Appendix E Ablation Study
E.1 Result Fusion
@@
E.2 Varying the Base Models
In this section, we use four different lightweight LLMs as the base models for CascadeNER, i.e., two versions of Qwen2 and Gemma with different parameters, namely Qwen2-1.5B, Qwen2-7B, Gemma-2B, and Gemma-7B. Gemma is another prominent lightweight LLM series, proposed by Google Team et al. (2024). These models represent the current best-performing lightweight LLMs. We use the dynamic categorized version of DynamicNER to fine-tune the instruct versions of the four models. Each model is fine-tuned separately on the corresponding dataset to obtain both an extractor and a classifier. The performance comparison of these combinations on a selected part of DynamicNER is shown in Table 9. Based on these results, Qwen2.5 outperform Gemma in multilingual tasks overall. Therefore, we choose Qwen2.5 as the base models for other experiments.
English Chinese Spanish Japanese
Qwen2.5-1.5B 62.8 58.9 55.7 54.1
Qwen2.5-7B 68.2 64.5 61.5 60.8
Gemma-2B 58.9 49.5 53.1 48.4
Gemma-7B 61.7 53.9 55.3 52.1
Table 9: F1 scores for CascadeNER with different base models on a selected part of DynamicNER
E.3 Context in Classification
@@
Appendix F LLM-based Methods Comparsion
@@
Appendix G Computational Resource Usage Record
@@
Appendix H Data Contamination Statement
@@
Appendix I Ethical Statement of DynamicNER
@@
For the annotators, each language in DynamicNER is annotated by two junior or higher-level students from the corresponding language departments at our university. Due to the double-blind review process, the annotators’ identities cannot be disclosed in this version. Each annotator receives extensive training and follows DynamicNER’s multi-granularity classification system to ensure consistent and accurate entity annotations across various languages and domains. The annotation process for each language are divided into two parts equally, with each annotator independently handling one part. After the initial annotation, the annotators revise their work based on the review results. For ambiguous terms or specialized domain terms, the annotators either collaborate with each other or consult experts to ensure the accuracy and reliability of the annotations.
In our writing, we use CahtGPT-4o Achiam et al. (2023) and Claude 3.5 Anthropic (2023) for assistance.
Appendix J Detailed Categories of DynamicNER
J.1 Person
@@
Appendix K More Categorization Quality Evaluation

### v5

@@ Appendix D Detail of DynamicNER
Although DynamicNER’s data volume is significantly smaller than existing multilingual NER datasets, it aligns with our primary goal: evaluating LLM-based NER in fine-grained, low-resource, and dynamic scenarios. In such contexts, data volume may not be decisive. We acknowledge that smaller data volumes may lead to higher variance and overfitting. To show the robustness of our work, we conduct additional experiments on existing datasets with the model finetuned on DynamicNER. As shown in Appendix E, especially Table 7 & 8, CascadeNER, trained on DynamicNER, demonstrates strong few-shot performance on other established benchmarks. This indicates that the models have learned robust, generalizable features, avoiding overfitting.
Appendix E More Experiment about CascadeNER
E.1 CascadeNER Setting
In the experiments of this section, CascadeNER always employs two Qwen2.5-7B base models, which are fine-tuned separately based on the corresponding part of the dynamic version of DynamicNER to obtain an extractor and a classifier. Potential data contamination about the fine-tuning is discussed in Appendix N. We evaluate CascadeNER’s performance in both few-shot and zero-shot scenarios, comparing it with supervised SOTAs and LLM-based baselines. For few-shot scenarios, the number of few-shot demonstrations is set to $3$, the same as the experiments on DynamicNER.
E.2 Baselines
@@
E.3 Dataset
@@
Few-shot Data Sampling.
Dataset Selection.
CoNLL2003.
@@
MultiCoNER & PAN-X.
MultiCoNER and PAN-X are two widely used multilingual datasets. MultiCoNER covers 6 entity types and 11 languages, while PAN-X includes 3 entity types and 282 languages. We use 6 and 8 overlapping languages from MultiCoNER and PAN-X with DynamicNER to evaluate CascadeNER’s multilingual capabilities. Notably, for the purpose of controlling variables, all methods requiring training are trained using multilingual joint training.
E.4 Experimental Results
As shown in Table 7 and 8, the results indicate that in low-resource scenarios, LLM-based methods achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music and FewNERD, and PAN-X and MultiCoNER in some languages, achieving new SOTA performance and highlighting its exceptional generalization and capability to handle complex entity categorization. However, when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X, indicating that supervised methods are still useful in some scenarios.
Appendix F Ablation Study
F.1 Result Fusion
@@
F.2 Context in Classification
@@
Model Base-Supervised Dynamic-Supervised Dynamic-Fewshot
en zh en zh en zh
G-1.5B 50.2 $\pm$ 1.3 44.1 $\pm$ 1.3 47.6 $\pm$ 1.5 41.2 $\pm$ 1.4 36.9 $\pm$ 1.6 35.8 $\pm$ 1.6
G-7B 55.1 $\pm$ 1.2 50.9 $\pm$ 1.2 52.3 $\pm$ 1.3 48.1 $\pm$ 1.3 42.7 $\pm$ 1.4 41.1 $\pm$ 1.4
G-GPT 62.4 $\pm$ 1.1 60.4 $\pm$ 1.2 60.6 $\pm$ 1.3 58.4 $\pm$ 1.3 49.2 $\pm$ 1.4 48.9 $\pm$ 1.3
P-1.5B 21.6 $\pm$ 1.6 20.7 $\pm$ 1.5 23.2 $\pm$ 1.7 22.7 $\pm$ 1.7 20.5 $\pm$ 2.0 19.9 $\pm$ 1.8
P-7B 41.1 $\pm$ 1.4 47.4 $\pm$ 1.4 44.3 $\pm$ 1.5 40.4 $\pm$ 1.5 39.8 $\pm$ 1.8 37.8 $\pm$ 1.8
P-GPT 49.7 $\pm$ 1.3 48.6 $\pm$ 1.3 53.0 $\pm$ 1.5 52.3 $\pm$ 1.4 49.4 $\pm$ 1.8 47.4 $\pm$ 1.9
C-1.5B 67.6 $\pm$ 1.1 64.0 $\pm$ 1.1 62.8 $\pm$ 1.2 58.9 $\pm$ 1.1 49.7 $\pm$ 1.3 48.5 $\pm$ 1.4
C-7B 73.8 $\pm$ 1.0 69.8 $\pm$ 1.0 68.2 $\pm$ 1.1 64.5 $\pm$ 1.1 55.7 $\pm$ 1.2 52.9 $\pm$ 1.2
C-GPT 77.1 $\pm$ 0.9 74.3 $\pm$ 0.9 73.1 $\pm$ 1.0 68.3 $\pm$ 1.0 61.3 $\pm$ 1.1 59.7 $\pm$ 1.1
Table 11: F1 Scores with 95% CI for English and Chinese. This table directly validates the statistical significance of our findings for representative languages. The consistently narrow CIs for CascadeNER highlight its high performance stability.
Appendix G Statistical Significance Analysis
To prove the statistical significance of our results on DynamicNER, we conduct an analysis to calculate 95% confidence intervals (CI) for our primary findings. This analysis validates whether the observed performance differences between methods, particularly the advantages of CascadeNER, are statistically meaningful.
We employ bootstrap resampling over the test set to compute the 95% CI for the F1 scores. For a test set of size $N$, we create a bootstrap sample by randomly drawing $N$ instances with replacement. The F1 score is then calculated on this new sample. By repeating this process 500 times, we obtain an empirical distribution of F1 scores. The 95% confidence interval is then determined by taking the 2.5th and 97.5th percentiles of this distribution. This analysis is performed on English and Chinese as two representative languages. The results are presented in Table 11.
The results from our statistical analysis yield two key insights. First, the performance gains of CascadeNER over the baseline methods are confirmed to be statistically significant, as the confidence intervals show clear separation in most cases, particularly for the stronger GPT-4o backbone. Second, CascadeNER consistently exhibits the narrowest confidence intervals across almost all settings. This is especially evident when compared to PromptNER, whose complex prompts lead to higher performance variance. This indicates that by decomposing the NER task into two simpler sub-tasks, CascadeNER not only achieves higher accuracy but also offers a more stable and reliable performance.
Appendix H LLM-based Methods Comparsion
@@
Appendix I Computational Resource Usage Record
@@
Appendix J Human Annotation Process
J.1 Manual Annotation
For the annotators, each language in DynamicNER is annotated by two junior or higher-level students from the corresponding language departments at our universities. Each annotator receives specific training and follows DynamicNER’s multi-granularity classification system to ensure consistent and accurate entity annotations across various languages and domains. The annotation process for each language are divided into two parts equally, with each annotator independently handling one part. After the initial annotation, the annotators revise their work based on the review results with the aid of Claude 3.5 Sonnet Anthropic (2023). For ambiguous terms or specialized domain terms, the annotators either collaborate with each other or consult experts via personal contact to ensure the accuracy and reliability of the annotation process.
Appendix K Inter-Annotator Agreement Analysis
To empirically validate the quality and semantic correctness of DynamicNER, we conduct a Inter-Annotator Agreement (IAA) analysis. We focus on the Dynamic Version, to ensure that the entity labels remain semantically robust and reliable even after undergoing the automated, metric-driven dynamic categorization process.
We select Cohen’s Kappa ($\kappa$) as our evaluation metric, as it is a widely recognized statistical measure for evaluating the reliability of judgments between two independent annotators. The process includes two stages. Firstly, we apply our stratified sampling method (detailed in Appendix E) to randomly sample 200 sentences each for English, French, Japanese, and Chinese from the Dynamic Version. These languages are chosen as representative cases for the analysis. The sampled sentences are then given to two annotators who are not involved in the original dataset creation. They perform the annotation in a double-blind setting, where neither annotator was aware of the other’s decisions or the original labels. The results of our IAA analysis are presented in Table 13.
Table 13: Cohen’s Kappa ($\kappa$) scores from the IAA analysis conducted on the dynamic version of DynamicNER.
Language Cohen’s Kappa ($\kappa$)
English 0.86
French 0.82
Chinese 0.83
Japanese 0.81
According to established interpretations of the Kappa statistic, the obtained scores indicate substantial agreement. This strong quantitative evidence, derived from a rigorous human-centric audit, directly confirms that the entity labels in our dataset are semantically robust and reliable. It serves as a crucial validation that our automated quality-control pipeline successfully preserves high-quality annotations, thereby addressing the concern about the semantic correctness of the dynamically generated labels.
Appendix L Finetuning Hyperparameter
To ensure the reproducibility of our experiments, this section provides a detailed record of the hyperparameter configurations used for fine-tuning the lightweight LLMs in CascadeNER.
Our implementation is based on the well-regarded MS-SWIFT framework Team (2024), and we employ Low-Rank Adaptation (LoRa) Hu et al. (2022) for Parameter-Efficient Fine-Tuning (PEFT) Houlsby et al. (2019). All fine-tuning and inference experiments are conducted on a server equipped with 8 NVIDIA RTX 3090 GPUs with 24GB VRAM for each GPU.
Hyperparameter Value
LoRA Parameters
Target Modules ALL
Rank (r) 8
Alpha ($\alpha$) 32
Dropout 0.05
Training Parameters
Learning Rate 1e-4
Optimizer AdamW
Weight Decay 0.01
Max Sequence Length 4096
Training Epochs Max 50
  - Best Epoch (1.5B model) 25
  - Best Epoch (7B model) 20
Effective Batch Size 32
Table 14: Hyperparameters used for fine-tuning CascadeNER models. The ’ALL’ setting for Target Modules indicates that LoRA was automatically applied to all suitable linear layers within the model.
The specific parameters for LoRA and the training process are detailed in Table 14. We maintain a consistent effective batch size of 32 for both models to ensure a fair comparison. For the 1.5B model, we use 4 GPUs with a per-device batch size of 4 and 2 gradient accumulation steps. For the 7B model, we use 8 GPUs with a per-device batch size of 1 and 4 gradient accumulation steps. We train each model for a maximum of 50 epochs, saving a checkpoint every 5 epochs and using an early stopping strategy based on the validation set performance. The best-performing checkpoints are then selected for the final evaluation.
Appendix M Prompt Example for CascadeNER
In this section, we present the examples of prompts used in CascadeNER to provide an intuitive demonstration of the method and to facilitate reproducibility. The example is provided in Figure 11. We provide complete prompts and corresponding json files for training in our github repository.
Figure 11: Examples for the complete prompts of CascadeNER.
Appendix N Data Contamination Statement
@@
Appendix O Ethical Statement
@@
In our writing, we use ChatGPT-4o Achiam et al. (2023) and Claude 3.5 Sonnet Anthropic (2023) for assistance.
Appendix P Detailed Categories of DynamicNER
P.1 Person
@@
Appendix Q More Categorization Quality Evaluation